please solve my problem: 2v+18=16-4(v+7)
Answer by checkley77(12844)

2v+18=16-4(v+7)
2v+18=16-4v-28
2v+4v=16-28-18
6v=-30
v=-30/6
v=-5 ans.
Proof:
2*-5+18=16-4(-5+7)
-10+18=16-4*2
8=16-8
8=8